A machine with a mass of 4kg fires a 45g bullet at a speed of 100m/s...
What is the recoil speed of the machine gun
1
Expert's answer
2020-03-27T10:51:19-0400

"M" =4 kg - mass of machine

"m" = 45 g - mass of bullet

"v" = 100 m/s - speed of bullet

"v'" - recoil speed of machine gun

Law of conservation of momentum:

"0=mv-Mv'"

Then "v'=mv\/M=0.045*100\/4=1.125" m/s